A mixture of 14% disinfectant solution is to be made from 10% and 20% disinfected solutions. How much of each solution should be used if 20 gallons of 14% solutions are needed?
Found 2 solutions by mananth, modi.kk91:
Answer by mananth(16946)

A mixture of 14% disinfectant solution is to be made from 10% and 20% disinfected solutions. How much of each solution should be used if 20 gallons of 14% solutions are needed?
10% solution be x gallons
20% solution will be 20-x gallons
..
Total 14% solution 20 gallons
..
Concentration eauation
0.1x+0.2(20-x)=20*0.14
0.1x+4-0.2x=2.8
-0.1x= -1.2
x= -1.2/ -0.1
x= 12 gallons of 10% solution
Balance 8 gallons will be 20% solution

Answer by modi.kk91(20)

by allegation rule
10% 20%


14%



20-14=6 and 14-10=4



ratio=6:4=3:2=========3x+2x=5x
ans= 3*4=12
2*4=8
5x=20
x=4